What is the formula for copper(II) phosphate? Capitalization and punctuation count.

formula:

How many atoms of phosphorus are in 9.309.30 mol of copper(II) phosphate?

number of P atoms =

please look at the question for accurate question
Answer #1

1)

Copper(II) ion is Cu2+

Phosphate anion is PO43-

The compound is:

Cu3(PO4)2

Answer: Cu3(PO4)2

2)

use:

number of molecules = number of mol * Avogadro’s number

number of molecules = 9.30 * 6.022*10^23 molecules

number of molecules = 5.60*10^24 molecules

1 molecule of Cu3(PO4)2 has 2 atoms of P

So,

Number of atoms of P = 2*number of molecules

= 2 * 5.60*10^24

= 1.12*10^25 atoms

Answer: 1.12*10^25 atoms